### v2

# Trustworthy Enhanced Multi-view Multi-modal Alzheimer’s Disease Prediction with Brain-wide Imaging Transcriptomics Data

Abstract
Brain transcriptomics provides insights into the molecular mechanisms by which the brain coordinates its functions and processes. However, existing multimodal methods for predicting Alzheimer’s disease (AD) primarily rely on imaging and sometimes genetic data, often neglecting the transcriptomic basis of brain. Furthermore, while striving to integrate complementary information between modalities, most studies overlook the informativeness disparities between modalities. Here, we propose TMM, a trusted multiview multimodal graph attention framework for AD diagnosis, using extensive brain-wide transcriptomics and imaging data. First, we construct view-specific brain regional co-function networks (RRIs) from transcriptomics and multimodal radiomics data to incorporate interaction information from both biomolecular and imaging perspectives. Next, we apply graph attention (GAT) processing to each RRI network to produce graph embeddings and employ cross-modal attention to fuse transcriptomics-derived embedding with each imaging-derived embedding. Finally, a novel true-false-harmonized class probability (TFCP) strategy is designed to assess and adaptively adjust the prediction confidence of each modality for AD diagnosis. We evaluate TMM using the AHBA database with brain-wide transcriptomics data and the ADNI database with three imaging modalities (AV45-PET, FDG-PET, and VBM-MRI). The results demonstrate the superiority of our method in identifying AD, EMCI, and LMCI compared to state-of-the-arts. Code and data are available at https://github.com/Yaolab-fantastic/xx.
Index Terms:
Imaging transcriptomics, trustworthy learning, graph attention, cross-modal attention, Alzheimer’s disease
I Introduction
Alzheimer’s disease (AD) is a progressive neurodegenerative disorder that predominantly affects the elderly, characterized by the gradual deterioration of cognitive functions, such as memory, reasoning, and decision-making [gaubert2021decision]. Studies have demonstrated that the pathological changes associated with AD begin to manifest years or even decades before the earliest clinical symptoms are observed [yao2019bioinfo].
Recent advances in acquiring biomedical data, coupled with the rapid advancements in machine learning and deep learning techniques, have significantly enhanced the development of computer-aided diagnosis of AD. The multifactorial nature of AD highlights the need for multi-modal learning, where numerous algorithms have been developed to capture and integrate complementary information from various modalities. Existing multimodal methods for AD prediction primarily focus on using brain imaging data. https://doi.org/10.1016/j.media.2022.102698;DOI:10.1109/TMI.2024.3386937 Although imaging modalities effectively quantify the structural and functional alterations in the brain, they do not account for the molecular mechanisms underlying these changes. A few studies have also incorporated genotyping data into the prediction models to enhance their performance. ref: https://link.springer.com/chapter/10.1007/978-3-031-43895-0_64,https://doi.org/10.1016/j.media.2023.102883,10.1109/TMI.2022.3162870 For example, Bi et al. constructed brain region-gene networks using fMRI and SNP data and aggregated the multimodal information through a modified generative adversarial network [bi2022novel]. Venugopalan et al. [venugopalan2021multimodal] employed deep learning to integrally analyze MRI, SNP, and clinical test data, using stacked denoising auto-encoders and 3D-CNNs to enhance AD, MCI, and CN classification. While genotypic data can characterize individuals, this upstream information does not directly explain cellular function and pathological states. Compared to genetics, downstream transcriptomics provides more direct insights by reflecting the actual activity and expression of genes within specific tissue contexts. This motivates us to incorporate brain-specific transcriptomics with imaging data into AD prediction.
Besides improving representation learning, there has been a growing emphasis on developing various fusion strategies to capture the complementary information present across modalities . For example, Song et al. [song2021graph] developed functional and structural graphs using fMRI and diffusion tensor imaging data, and combined the multimodal information into edges through an efficient calibration mechanism. Bi et al. [bi2023community] developed a community graph convolutional neural network to model brain region-gene interactions, using an affinity aggregation model to enhance interpretability and classification performance in AD diagnosis.Although effective, existing multimodal algorithms often lack a dynamical perception of the informativeness of each modality for different samples, which could otherwise enhance the trustworthiness, stability, and explainability of these methods. Evidence has shown that the informativeness of a modality typically varies across different samples [Dynamicsref[28, 49]], which motivates efforts to model modality informativeness to enhance fusion effectiveness .
To address the aforementioned issues, we propose a trustworthy enhanced multi-view multi-modal graph network (TMM) that integrates brain-wide transcriptomics and multi-modal radiomics data for predicting AD in an adaptive and trustworthy manner. In this context, multi-view refers to the incorporation of both transcriptomics and radiomics data, while multi-modal pertains to including different imaging modalities. As illustrated in Figure LABEL:fig1_overview, we begin by constructing view-specific region of interest (ROI) co-function networks (RRIs) from brain-wide gene expression data and brain imaging data. These networks capture the molecular, structural, and functional relationships underlying different brain regions. Graph attention networks (GAT) are then applied to generate sample-wise embeddings, followed by cross-modal attention to fuse representations from different views. In the fusion stage, we propose a novel true-false-harmonized class probability (TFCP) mechanism to measure modal confidence, facilitating the adaptive perception of and response to variations in modality informativeness. Our main contributions can be summarized as follows:
  • We propose a multi-view, multi-modal learning framework for AD prediction that integrates region-specific transcriptomics and brain imaging data. This approach models the co-functional relationships of ROIs by combining insights from upstream biomolecular activities and downstream anatomical and functional characteristics.
  • We propose a novel modal confidence learning strategy that estimates the harmonized true and false class probability to dynamically assess the modality informativeness. This design enables a more robust and trustworthy multimodal prediction.
  • Our proposed model significantly outperforms state-of-the-art methods in predicting AD and MCI. A series of ablation studies robustly validate the effectiveness of the proposed TMM framework. Furthermore, we identify important ROIs and demonstrate their functional roles in brain cognition.
II Related work
Multimodal AD prediction. Existing multimodal methods for AD typically utilize brain imaging data (e.g., MRI, PET) because of their easy accessibility and early diagnosis potential. Common deep-learning architectures have been widely applied to learn embeddings of each individual modality, for example, convolutional neural network (CNN) , recurrent neural network (RNN) , autoencoder (AE) , generative adversarial network (GAN)  and transformer . Most recently, graph neural networks (GNNs) and their variants have gained significant attention thanks to their natural advantages in capturing interaction patterns between ROIs and their adaptability to different data types .
Besides imaging, genetics data have also been incorporated to enhance prediction performance [e.g., Xia-anBi, LiShen&LifangHe]; however, genetic variants do not directly contribute to the pathogenesis of AD but instead influence the disease through downstream functional elements (e.g., gene expression). Although high-throughput sequencing techniques have greatly accelerated the acquisition of omics data, their application in brain disorders is challenging due to the invasive nature of sample collection. The scarcity of sample-level brain transcriptomics data inspires us to explore using brain-wide transcriptomics data from healthy individuals as prior knowledge for multimodal AD prediction.
Confidence learning. Although various integration strategies have been proposed (e.g., early , intermediate , late , and hybrid  fusion), adapting to modal discrepancies remains challenging, particularly for heterogeneous and imbalanced multimodal data. By assessing and incorporating the predictive uncertainty of each modality into the fusion process, confidence learning has shown significant performance in multimodal prediction. Bayesian approaches introduce probability distributions over the model parameters to estimate uncertainty . Ensemble methods train multiple models independently and use prediction variance as uncertainty . Dropout-based methods, such as MC Dropout , leverage dropout regularization to obtain multiple predictions. Confidence calibration methods  aim to align the predicted probabilities with the true probabilities of correctness, thereby improving the reliability of uncertainty estimates. Han et al. [Han_2022_CVPR] introduced the true class probability criterion to assess the uncertainty of each modality. Luo et al. [luo2024teminet] estimated the parameters of the Dirichlet distribution based on the prediction of each modality. Zheng et al. [Zheng2023multilevel] assess the decision confidence of each modality by constraining the orthogonalities between each pair of predicted classes. In this paper, we designed a novel confidence calibration criterion to help estimate the informativeness of each modality for trustworthy integration.
Figure 1: Framework overview. (a) Transcriptomics-specific RRI network (T-RRI) is constructed from the brain-wide gene expression data, and radiomics-specific RRI network is derived from each imaging data. (b) Three imaging modalities including VBM, FDG, and AV45 are employed in our analysis. (c) The TMM architecture. (i) For each sample, modality-specific T-RRI and R-RRI networks are constructed by mapping ROI measurements to each node. Multi-level GAT is applied to capture interactions within ROIs for each view, generating view-specific representations, followed by cross-view attention to create multi-view embeddings. (ii) The true-false-harmonized class probability (TFCP) is designed to estimate the prediction confidence for each modality. (iii) Cross-modal attention mechanisms fuse multiple modalities to deliver the final predictive output. (d-f) illustrates the details of cross-view graph fusion, the TFCP mechanism, and the cross-attention mechanisms, respectively. (g) The identified ROI biomarkers are annotated for their functional co-activation.
III Methods
In the following sections, we denote matrices as boldface uppercase letters and vectors as boldface lowercase ones. Let $\mathbf{T}=\{\mathbf{t}_{1},\dots,\mathbf{t}_{d}\}\in\mathbb{R}^{n_{g}\times d}$ represent the brain-wide transcriptomics data, with $n_{g}$ denoting the number of genes and $d$ denoting the number of ROIs. Let $m\in[1,\dots,M]$ denotes the $m$-th imaging modality, $\mathbf{X}^{m}=\{\mathbf{x}^{m}_{1},\dots,\mathbf{x}^{m}_{d}\}\in\mathbb{R}^{n\times d}$ represents the $m$-th imaging feature matrix, and $\mathbf{y}=[y_{1},\dots,y_{n}]\in\mathbb{R}^{n}$ denotes the label vector, where $n$ is the number of samples and $d$ is the number of ROIs.
We now detail the proposed TMM framework. Initially, one transcriptomics-specific network (T-RRI) and $M$ radiomics-specific RRI networks (R-RRIs) are constructed from brain-wide gene expression data and each imaging modality dataset, respectively. For each imaging modality, ROI-level imaging measurements are assigned as node features to both the T-RRI and R-RRI networks, generating graph representations from transcriptomic and imaging views. Graph attention networks are then applied to each RRI network to generate view-specific embeddings, followed by cross-modal attention mechanisms to fuse multi-view representations. Subsequently, a confidence learning network is specifically designed to generate more reliable embeddings, with cross-modal attention employed to fuse these multimodal representations for final prediction. Below, we delve into the RRI co-functional networks construction, multi-view fusion, confidence learning module, and the multi-modal integration, detailed in Sections III-A,  III-B and  III-C, respectively. The overall architecture and important modules are illustrated in Figure 1.
III-A Sample-wise Co-functional Network Construction
Two types of RRI networks (T-RRI and R-RRI) are constructed from transcriptomics and imaging data, respectively, to model the co-functionality of ROIs and facilitate graph-based network implementation. Given the brain-wide gene expression matrix $\mathbf{T}\in\mathbb{R}^{n_{g}\times d}$, we derive the edge matrix $\mathbf{E}^{t}\in\mathbb{R}^{d\times d}$ by thresholding the adjacency matrix, which measures the correlation coefficient for each pair of ROIs across all expressed genes:
$\text{E}^{t}_{ij}=\left\{\begin{array}[]{cl}1&\text{ if }r(\mathbf{t}_{i},\mathbf{t}_{j})\geq\lambda\\ 0&\text{ otherwise}\end{array}\right.,$ (1)
where $r(\mathbf{t}_{i},\mathbf{t}_{j})=\frac{\sum_{k=1}^{g}(\mathbf{t}_{ik}-\overline{\mathbf{t}_{i}})\cdot(\mathbf{t}_{jk}-\overline{\mathbf{t}_{j}})}{\sqrt{\sum_{k=1}^{g}(\mathbf{t}_{ik}-\overline{\mathbf{t}_{i}})^{2}}\cdot\sqrt{\sum_{k=1}^{g}(\mathbf{t}_{jk}-\overline{\mathbf{t}_{j}})^{2}}}$ computes the Pearson correlation coefficient (PCC) between the expressions of ROIs $\mathbf{t}_{i}$ and $\mathbf{t}_{j}$, and $\lambda$ is a hyperparameter for thresholding.
We apply the same strategy to each imaging modality, obtaining $M$ imaging-specific edge matrix $\mathbf{E}^{m}\in\mathbb{R}^{d\times d}$ (where $m\in[1,\dots,M]$), with each element $\text{E}^{m}_{i,j}$ representing the PCC between imaging measures of each ROI pair across all $n$ samples. Across experiments, we set the thresholds $\lambda_{t}=0.2$ for transcriptomics-based networks and $\lambda_{r}=0.1$ for radiomics-based networks, respectively.
Afterward, for each imaging modality, we devise one T-RRI network and one R-RRI network for each subject by assigning ROI-level imaging measurements to their corresponding nodes as features. Consequently, each subject is associated with $M$ T-RRI graphs and $M$ R-RRI graphs. Notably, the $M$ T-RRI graphs share the same edge matrix but differ in node features, while the $M$ R-RRI graphs each have unique edge matrices and distinct node features.
III-B Multi-view Fusion for Modal-Specific Representation
View-specific representation. Given their ability to effectively aggregate neighboring relationships within graph data, we utilize GATs for multi-view sample representation learning. Specifically, multi-level GATs are applied to each T-RRI and R-RRI graph, enabling the sample representations from both transcriptomic and different imaging modal views. Illustrated with the $m$-th modality, we take $\mathcal{G}^{m}_{t,0}=(\mathbf{X}^{m},\mathbf{E}^{t})$ and $\mathcal{G}^{m}_{r,0}=(\mathbf{X}^{m},\mathbf{E}^{m})$ as input T-RRI and R-RRI graphs and stack graph attention layers with multi-head attention to build the GAT for each graph. Each layer is defined as:
$\boldsymbol{h}_{u}^{\prime}=\|_{k=1}^{K}\sigma\left(\sum\limits_{v\in{\mathcal{N}_{u}}}\alpha_{uv}^{k}\mathbf{W}^{k}\boldsymbol{h}_{v}\right),$ (2)
where $\parallel$ is concatenation of $K$ heads, $\boldsymbol{h}_{v}$ is the input features of node $v$, $\alpha_{uv}^{k}$ is the $k$-th normalized attention coefficients, $\mathbf{W}^{k}$ is the weight matrix of head $k$, $\mathcal{N}_{u}$ is the first-order neighbors of node $u$, and $\sigma(\cdot)$ is a nonlinear activation function. $\alpha_{uv}$ is calculated from the attention mechanism $\alpha_{u,v}=\frac{\exp(a(\boldsymbol{h}_{u},\boldsymbol{h}_{v}))}{\sum\limits_{v^{\prime}\in\mathcal{N}_{u}}\exp(a(\boldsymbol{h}_{u},\boldsymbol{h}_{v^{\prime}}))}$, with $a(\boldsymbol{h}_{u},\boldsymbol{h}_{v})$ representing the importance of $v$ to $u$.
We further adopted a multi-level strategy to enhance the aggregation of information within brain circuits. Specifically, higher-level graphs $\mathcal{G}^{m}_{t,1}$ and $\mathcal{G}^{m}_{r,1}$ are generated by applying the GAT processes to initial graphs $\mathcal{G}^{m}_{t,0}$ and $\mathcal{G}^{m}_{r,0}$. Then, we derived the next level graphs, $\mathcal{G}^{m}_{t,2}$ and $\mathcal{G}^{m}_{r,2}$ from $\mathcal{G}^{m}_{t,1}$ and $\mathcal{G}^{m}_{r,1}$ respectively. For each modality, graph embeddings from all three levels are concatenated to produce view-specific representations, denoted as $\mathbf{F}^{m}_{t}$ and $\mathbf{F}^{m}_{r}$.
Multi-view fusion. Within each modality, we introduce cross-view attention mechanisms to integrate representations from the T-RRI and R-RRI networks, enabling them to complement and enhance each other:
$\mathbf{Z}^{m}_{t}$ $=\text{Att}\Big{(}\mathbf{F}^{m}_{t}\mathbf{W}^{Q}_{t},\mathbf{F}^{m}_{r}\mathbf{W}^{K}_{r},\mathbf{F}^{m}_{r}\mathbf{W}^{V}_{r}\Big{)},$ (3)
$\mathbf{Z}^{m}_{r}$ $=\text{Att}\Big{(}\mathbf{F}^{m}_{r}\mathbf{W}^{Q}_{r},\mathbf{F}^{m}_{t}\mathbf{W}^{K}_{t},\mathbf{F}^{m}_{t}\mathbf{W}^{V}_{t}\Big{)},$
$\mathbf{Z}^{m}$ $=\mathbf{Z}^{m}_{t}\operatorname*{\mathchoice{\Big{\|}}{\big{\|}}{\|}{\|}}\mathbf{Z}^{m}_{r},$
[…]
In addition to fusing transcriptomics and radiomics information, we also train a modal-specific classifier to incorporate within-modal information into prediction. In particular, for the $m$-th modality, a GAT classifier is trained as:
$\mathcal{L}_{\text{GAT}}^{m}=\sum\limits_{i=1}^{n}\mathcal{L}_{\text{CE}}(\hat{\mathbf{y}}_{i}^{m},\mathbf{y}_{i}),$ (4)
where $\mathcal{L}_{\text{CE}}$ is the cross-entropy loss, ${n}$ is the number of training samples, $\mathbf{y}_{i}$ is the true label, and $\hat{\mathbf{y}}_{i}^{m}$ is the prediction from $\mathbf{Z}^{m}$.
III-C Modality-level Confidence: True-False-Harmonized Class Probability
In multimodal learning, data heterogeneity is a common challenge, with the discriminative capability of each modality varying across different samples. Therefore, it is crucial to develop algorithms that can adaptively perceive and respond to these variations in informativeness. By transforming the assessment of modal informativeness into an evaluation of the confidence levels associated with modal classification performance, we can effectively estimate the modality informativeness through its prediction confidence. Here, we propose a novel true-false-harmonized class probability (TFCP) to approximate the prediction confidence of each modality.
The confidence criterion measures how confident the model is in its predictions by correlating high prediction certainty with greater values and vice versa. For a classifier $f:\mathbf{X}\rightarrow\mathbf{y}$, the traditional confidence criterion utilizes the maximum class probability (MCP): $\text{MCP}(\mathbf{x}_{i})=P(\widehat{y_{i}}|\mathbf{x}_{i})$, where $\widehat{y_{i}}$ is the class with the largest softmax probability. However, it can be observed that this method assigns high confidence (i.e., high softmax probability) to both correct and incorrect predictions, resulting in overconfidence for failure predictions. True class probability (TCP) criterion [corbiere2019addressing] has been proposed to calibrate incorrect predictions. Instead of assigning the maximum class probability, TCP uses the softmax probability of the true class $y_{i}^{*}$ as prediction confidence: $\text{TCP}(\mathbf{x}_{i})=P(y_{i}^{*}|\mathbf{x}_{i})$, where $y_{i}^{*}$ is the class with the largest softmax probability. The TCP strategy exclusively models the certainty of the true class while neglecting the uncertainty of the untrue classes, which could result in a biased and unstable confidence approximation.
To this end, we propose considering both true and false class probabilities to aggregate evidence from these two perspectives. Specifically, we design a harmonized criterion, named $\text{Conf}_{\text{TFCP}}$, for the prediction confidence:
$\text{TFCP}(\mathbf{x}_{i})$ $=\frac{2}{1/{\text{TCP}(\mathbf{x}_{i})}+1/{1-\text{FCP}(\mathbf{x}_{i})}}$ (5)
$=\frac{2}{1/{P(y_{i}^{*}|\mathbf{x}_{i})}+1/{(1-P(\overline{y_{i}^{*}}|\mathbf{x}_{i}))}},$
where $\text{FCP}(\mathbf{x}_{i})=P(\overline{y_{i}^{*}}|\mathbf{x}_{i})$ is the prediction probability of false (or untrue) class(es).
It should be noted that neither the TCP nor the FCP can be directly applied to test samples because they both require label information. To address this issue, we introduce two confidence networks to approximate the TCP and FCP, respectively. As shown in Figure 1(D), a TCP confidence network with parameters $\theta_{T}$ and an FCP confidence network with parameters $\theta_{F}$ are constructed based on the sample representations of each modality to generate certainty and uncertainty scores. These networks are trained to minimize the discrepancy between predicted and actual scores:
$\mathcal{L}_{\mathrm{Conf}}^{m}\!=\!\left\|\text{TFCP}(\mathbf{X}^{m})-\widehat{\text{TFCP}}(\mathbf{X}^{m},\theta^{m}_{T},\theta^{m}_{F})\right\|^{2}\!+\!\mathcal{L}_{\text{Cls}}^{m}.$ (6)
Here, both the TCP and FCP networks are built upon a classification network that is trained using the cross-entropy loss $\mathcal{L}_{\text{Cls}}^{m}$.
III-D Multi-modal Fusion: Cross-modal Attention
From the previous sections, multi-view sample representations for each modality have been derived, and modality informativeness has also been assessed, denoted as $\mathbf{Z}^{m}$ and $\widehat{\text{TFCP}}^{m}$, respectively. By incorporating the modality confidence into the corresponding representations, we can obtain trustworthy representation for each modality:
$\mathbf{H}^{m}=\widehat{\text{TFCP}}^{m}\cdot\mathbf{Z}^{m}$ (7)
Now, we employ cross-modal attention mechanisms to enhance each modality by leveraging insights from others, subsequently concatenating the enriched modal representations for final prediction:
$\mathbf{U}^{m}$ $=\|_{j=1,j\neq m}^{M}\text{Att}\Big{(}\mathbf{H}^{m}\mathbf{W}^{Q}_{m},\mathbf{H}^{j}\mathbf{W}^{K}_{j},\mathbf{H}^{j}\mathbf{W}^{V}_{j}\Big{)},$ (8)
$\mathbf{U}$ $=\|_{m=1}^{M}\mathbf{U}^{m}.$
III-E Objective optimization.
The overall loss is composed of the modality-specific classification loss (Eq. 4), the TFCP loss (Eq. 6), and the cross-entropy loss of the final classification:
$\mathcal{L}=\eta_{1}\sum\limits_{m=1}^{M}\mathcal{L}_{\text{GAT}}^{m}+\eta_{2}\sum\limits_{m=1}^{M}\mathcal{L}_{\text{Conf}}^{m}+{\mathcal{L}_{\text{Final}}},$ (9)
where $\eta_{1}$ and $\eta_{2}$ denote hyperparameters for adjusting different losses. We set $\eta_{1}=1$ and $\eta_{2}=1$ across our experiments.
IV Experiments and Results
IV-A Dataset
AHBA dataset. Brain-wide transcriptomics data are sourced from the Allen human brain atlas (AHBA, human.brain-map.org), including over 58k probes sampled across $3,702$ spatially distinct brain locations from six donors. We utilize the abagen toolbox [abagen] for data preprocessing and employ the AAL atlas to map brain locations to ROIs, resulting in a number of 15,633 genes expressed across 116 ROIs. To ensure the introduced prior knowledge closely aligns with the underlying pathogenic mechanisms of AD, we further filter the genes using large-scale AD meta-GWAS data from the International Genomics of Alzheimer’s Project (IGAP) [kunkle2019]. Specifically, we use the MAGMA tool to derive gene-level $p$-values from SNP-level GWAS results, keeping those with nominally significant $p$-value (i.e., $p<0.05$) into consideration. This process yields a total of 1,216 genes across 116 ROIs for constructing the T-RRI edge matrix.
TABLE I: Demographic characteristics of the ADNI subjects.
Subjects NC EMCI LMCI AD
Number 221 275 190 165
Sex(M/F) 114/107 155/120 111/79 99/66
Age 76.34±6.58 71.50±7.12 74.08±8.55 75.35±7.88
Education 16.40±2.66 16.09±2.61 16.36±2.79 15.87±2.72
ADNI dataset. We conduct experimental validation using subjects from the Alzheimer’s Disease Neuroimaging Initiative database (ADNI, https://adni.loni.usc.edu/). Multimodal imaging data–including AV45 PET, FDG PET, and VBM MRI, along with corresponding diagnosis labels (NC, EMCI, LMCI, and AD) for each visit, are collected from 851 ADNI participants. Each imaging scan is processed according to established protocols [yao2019bioinfo], and measurements for 116 ROIs are derived from each modality using the AAL atlas. Detailed characteristic information and the number of subjects in each diagnostic group are listed in Table I.
Benchmark methods. We compare the proposed TMM with nine multimodal competitors, including three single-modal classifiers with early fusion (SVM, Lasso, XGBoost, and fully connected NN), three models with intermediate/late concatenation (GRidge [van2016better], BSPLSDA [singh2019diablo], and gated multimodal fusion (GMU [arevalo2017gated])), and two methods with advanced representation and fusion designs (Mogonet [2021MOGONET] and Dynamics [Han2022MultimodalDD]).
TABLE II: The comparison results on four ADNI classification tasks.
Method NC vs. AD NC vs. LMCI NC vs. EMCI EMCI vs. LMCI
ACC F1 AUC ACC F1 AUC ACC F1 AUC ACC F1 AUC
SVM 90.2±2.3 89.8±1.2 94.1±4.2 65.4±2.3 63.5±2.5 72.1±3.3 67.7±1.2 78.5±1.4 64.9±1.7 62.4±3.1 69.3±3.0 65.5±3.0
Lasso 91.2±3.0 89.8±3.6 96.6±1.4 67.7±4.6 65.6±5.5 76.7±3.1 66.5±3.4 76.2±2.6 66.7±3.2 65.4±3.2 70.7±3.8 68.0±2.0
XGBoot 91.2±1.4 89.7±1.6 96.8±0.8 70.1±1.2 67.3±1.9 76.0±1.5 69.9±3.5 80.3±2.3 61.7±2.7 64.5±1.9 71.6±1.9 67.0±3.0
NN 88.9±2.2 87.5±2.5 51.9±7.0 69.1±2.6 67.6±3.3 52.1±6.3 69.7±2.1 79.4±1.9 60.9±2.8 64.8±4.8 71.0±3.7 47.7±3.7
GRridge 80.6±2.7 78.2±3.3 88.3±2.8 62.7±2.8 61.9±4.0 68.6±3.4 58.1±2.0 67.7±2.7 57.7±2.7 60.4±3.4 65.7±2.9 62.7±2.9
BSPLSDA 83.0±1.6 66.4±2.2 87.8±2.7 63.5±1.6 63.7±1.9 71.8±2.5 58.3±1.7 67.7±2.3 60.2±2.8 61.4±2.7 69.3±3.0 62.5±2.6
GMU 91.3±2.6 90.2±3.1 95.6±2.8 70.9±2.4 68.4±1.5 77.2±1.7 68.4±2.7 78.3±2.8 58.6±2.7 68.5±2.3 64.8±3.3 74.0±2.2
Mogonet 95.3±1.3 95.9±1.7 91.4±1.5 78.6±2.1 77.0±1.7 82.7±2.2 72.5±1.9 69.6±2.7 80.1±2.2 74.8±3.0 78.4±3.4 78.1±3.5
Dynamics 96.5±0.6 95.7±1.1 96.4±0.8 80.4±1.6 78.9±1.6 84.1±1.8 75.8±1.6 70.5±2.3 78.3±1.7 76.9±2.6 79.7±2.7 77.6±2.2
Ours 98.0±1.1* 97.7±1.3* 98.3±0.5* 83.7±1.2* 81.3±1.3* 85.9±1.6 78.6±1.5* 84.4±2.9* 77.3±1.2 80.0±2.0* 82.7±3.3 78.7±2.6
Best is in Bold. Suboptimal results are underlined. ‘*’ indicates that our TMM is significantly better (t-test $p<0.05$) than the suboptimal method.
Evaluation metrics. We employ accuracy (ACC), F1 score (F1), and the area under the receiver operating characteristic curve (AUC) to evaluate the performance of the methods. Five-fold cross-validation is performed to calculate the mean and standard deviation of the results. T-tests are performed to evaluate the significance of improvements achieved by our method over state-of-the-art methods.
Implementation details. We develop the model using PyTorch 2.1.0 and Adam as the optimizer. Specifically, we train the model for 2000 epochs with a learning rate (lr) of 1e-3 and weight decay (wd) of 1e-4. We implement all experiments on an RTX 4090 GPU with 24GB of memory.
IV-B Comparison with the state-of-the-art
Four tasks, including NC vs. AD, NC vs. EMCI, NC vs. LMCI, and EMCI vs. LMCI, are designed to validate the performance of our proposed method. Since the transcriptomics knowledge is not specific to ADNI subjects and cannot be directly incorporated, we are limited to using only multimodal imaging data for the comparative methods.
Table II shows the comparison results where our model illustrates superior performance across four tasks. Particularly, TMM achieves significant improvements in most evaluation metrics over the suboptimal results. This highlights the advantages of integrating prior transcriptomics knowledge through the well-designed multi-view graphs and the proposed trustworthy strategy.
IV-C Ablation studies
TABLE III: Ablation study of the key components in TMM.
T-RRI R-RRI NC vs. AD NC vs. LMCI NC vs. EMCI EMCI vs. LMCI
ACC F1 AUC ACC F1 AUC ACC F1 AUC ACC F1 AUC
98.0±1.1 97.7±1.3 98.3±0.5 83.7±1.2 81.3±1.3 85.9±1.6 78.6±1.5 84.4±2.9 77.3±1.2 80.1±1.7 82.4±2.6 78.5±2.3
97.2±1.2 96.3±1.8 98.0±1.2 82.0±1.3 79.6±2.1 83.4±2.7 76.4±2.2 82.1±3.2 75.6±2.5 77.8±2.5 80.6±2.9 76.3±2.8
96.7±1.4 95.9±2.1 97.2±1.4 81.4±1.6 78.2±1.4 83.6±1.8 77.3±1.4 82.8±2.7 76.8±1.6 78.3±2.3 81.8±2.4 77.5±2.0
95.6±2.6 94.8±2.2 96.3±2.7 79.8±2.4 76.3±2.9 81.7±2.6 75.4±2.3 80.6±2.5 75.4±1.7 76.3±2.7 79.3±3.2 76.7±2.3
Confidence NC vs. AD NC vs. LMCI NC vs. EMCI EMCI vs. LMCI
ACC F1 AUC ACC F1 AUC ACC F1 AUC ACC F1 AUC
TFCP 98.0±1.1 97.7±1.3 98.3±0.5 83.7±1.2 81.3±1.3 85.9±1.6 78.6±1.5 84.4±2.9 77.3±1.2 80.1±1.7 82.4±2.6 78.5±2.3
TCP 97.2±1.1 97.3±1.1 98.1±0.6 81.4±1.6 80.2±1.4 83.2±2.0 78.3±1.2 83.6±2.6 77.0±0.8 79.5±1.4 82.1±1.9 78.2±2.2
NN 96.6±1.6 96.0±2.0 97.3±1.6 79.5±2.3 78.5±2.6 81.5±2.7 77.4±1.5 82.2±2.7 76.3±1.4 77.4±2.4 80.1±2.8 75.2±2.6
Effectiveness of T-RRI, R-RRI and TFCP. We conduct ablation studies to evaluate the effectiveness of including the transcriptomics-derived RRI knowledge and the TFCP module. We also examine the informativeness of disease-specific R-RRI networks and the efficacy of cross-attention in information consolidation.
For comparative analysis, we accordingly remove the T-RRI network and three R-RRI networks, replace the TFCP with TCP and a fully connected neural network (NN) and replace the cross-view and cross-modal attentions with NN. We conduct ablation experiments on xxx tasks. Table III shows the results, from which we have the following observations: 1).
Figure 2: Performance comparison of different modality combinations on NC vs. AD classification.
IV-D Evaluations of modality contributions and parameter analysis
Modality contributions. To illustrate how multiple imaging modalities can offer complementary insights, we compared the classification performance using various combinations of imaging modalities. The results, displayed in Figure 2(b), reveal that: 1) different imaging modalities possess varying levels of discriminative capabilities, with VBM outperforming both FDG and AV45; 2) the integration of additional imaging modalities consistently yields better performance than using any subset alone, highlighting the unique contribution of each imaging modality and affirming the ability of our model in effectively modeling cross-modal information.
Hyperparameter analysis. We evaluate the hyperparameters for thresholding in the T-RRI and R-RRI networks. The parameters $\lambda_{t}$ and $\lambda_{r}$ are automatically tuned from the NC vs. AD task within the range of $[0.1,0.2,\cdots,0.6]$. Figure 3(a) shows the grid search results, based on which we select the optimal parameter combination of $\lambda_{t}=0.2$ and $\lambda_{r}=0.1$ for use across our experiments. We observe that variations in the thresholding parameters have minimal impact on performance, demonstrating the robustness of our proposed method.
Figure 3: Performance comparison of difference hyperparameters.
IV-E Identified biomarkers and functional annotation
For each classification task, feature ablation experiments [Dai_2021_WACV] are conducted to assess and rank the importance of ROIs. Table LABEL:tab:biomarker lists the top xx ROIs identified from each task.
V Conclusion
In this paper, we propose an adaptive multiomics integration framework introducing multi-view knowledge transfer for robust diagnosis classification. Graph contrastive pretraining is introduced to derive intra-omics patterns in an unsupervised manner. Cross-omics distillation facilitates the automatic transfer of inter-omics information from stronger to weaker modalities. Extensive experiments on benchmark datasets demonstrate the effectiveness of our approach. Concurrently, we have identified biomarkers that significantly impact disease inference, offering vital insights and pathways for elucidating pathological mechanisms and developing novel therapeutic approaches.